OpenStudy (anonymous):

I am to use the seven step method in my book to graph the following rational function of f(x)=x+2/x^2+x-6 HOWEVER my book doesn't have anything like this as an example with the seven steps. Just a loose idea of how to do it. Can anyone help?

OpenStudy (anonymous):

\[f(x)=\frac{ x+2 }{x ^{2}+x-6 } \] The problem asks you to determine the symmetry, find the y intercept, find the x intercept and the vertical asymptote

OpenStudy (campbell_st):

well there are vertical asymptotes at x= -3, x = 2 so there are 3 parts to the curve... it has a horizontal asymptote when x > 2 and x < -3 between the curves its continuous. then you have the x-intercept let x + 2 = 0 and solve for x.

OpenStudy (campbell_st):

vertical asymptotes from from the denomiantor what values of x make x^2 + x - 6 = 0 the values that do make it zero are the asymptotes

OpenStudy (anonymous):

Okay. :-) So if you were to type the equation of the asymptotes what would you have?

OpenStudy (campbell_st):

x = -3 and x = 2 here is an online graphing site that will help your understanding https://www.desmos.com/calculator

OpenStudy (anonymous):

I understand it better thank you. I have located the inter x and y intercepts but I am still unsure of the symmetry.